OpenStudy (anonymous):

PLEASE HELP! A 0.900 sample of toluene, C7H8, was completely burned a bomb calorimeter containing 4560.g of water which increased in temperature from 23.800C to 25.718C. What is Delta E for the reaction in kJ/mole C7H8? The heat capacity of the calorimeter was 780. J/C. The specific heat of water is 4.184 J/gC. The answer to this question is -3900 kJ/mol. If anyone can help me know how to solve this problem it could be greatly appreciated. You would be my lifesaver! Please Help! Thank you. :)

OpenStudy (aaronq):

You're finding the heat evolved by combustion by monitoring the water \(\sf \large \Delta H_{rxn}=m_{water}*C_p*(T_f-T_i)\) After, you need to scale it by taking into account the number of moles (n) of toluene used, to find the standard enthalpy of combustion (with units specified - note you will have to convert J to kJ at the end). \(\sf \large \Delta H_{rxn}=H^o*n_{toluene}\)

OpenStudy (anonymous):

Thank you so much! I asked my lab teacher, the tutor, and my chemistry I teacher to help me solve the problem, and none of them could help me solve the Chem II problem.

OpenStudy (aaronq):

You should try it out, if you can't with what i gave you, i'll do it for you

OpenStudy (anonymous):

Can you tell me what the mwater and the Cp stand for?

OpenStudy (aaronq):

\(m_{water}\)=mass of water \(C_p\) is the specific heat capacity

OpenStudy (anonymous):

So would the Cp be the heat capacity that =780 J/C?

OpenStudy (aaronq):

Oh i didn't read that, you need to account the heat absorbed by the calorimeter, so you need this extended formula: \(\Delta H_{rxn}=m_{water}*C_p*(T_f-T_i)+C_p*(T_f-T_i)\) water calorimeter

OpenStudy (anonymous):

Is the temperature of the calorimeter the same temperature of the water?

OpenStudy (aaronq):

yeah, they're in thermal equilibrium

OpenStudy (anonymous):

If you possibly can like take me step by step through it. I think I plugged in the right numbers. It should be Delta Hrzn = 49.57 mol (I went ahead and converted to moles) * 4.184*1.918*780*1.918 = 547094.7066. But now I don't know how to use the other equation.

OpenStudy (aaronq):

\(\Delta H_{rxn}\) should give you a value in terms of energy, not moles.

OpenStudy (aaronq):

I have to go, try it again and i'll check it when i'm on again

OpenStudy (anonymous):

Ok thank you.

OpenStudy (aaronq):

so it's like this: \(\sf \Delta H_{rxn}=(4560~g)* ( 4.184~ J/gC)*(25.718-23.800)C+780 J/C*(25.718-23.800)C\) (Note how the units cancel out and you're left with Joules) \(\sf \Delta H_{rxn}=38089.63872~J=38.09~ kJ\) Because the water and the calorimeter absorbed the heat (endothermic) this is a positive value.We need to change it to a negative value because the reaction is exothermic. --------------------------------------------------------- Then use the other formula to scale it up to 1 mole of toluene \(\sf-\Delta H_{rxn}=H^o_{combustion}*n_{toluene}\) \(\sf H^o_{combustionn}=\dfrac{-\Delta H_{rxn}}{n_{toluene}}=\dfrac{-38.09~ kJ}{0.9~mol}=-42.22 ~kJ/mol\) This number is a little different than the one you posted, but i'm 99.99% sure this is right.

OpenStudy (anonymous):

I'm sorry that I couldn't get back to you sooner, but thank you so much! You really helped, and you were right. The teacher said there was a mistake in the answer choices and the answer key on that question. Thank you again! :)

OpenStudy (aaronq):

Great, i'm glad you got it all figured out! no problem!